# Changes in Real Time: Online Scene Change Detection with Multi-View Fusion

### v2

@@ 3.3 Extracting Change Cues
Pixel-level change cues: We quantify the differences between $(I_{\text{inf}}^{k},I_{\text{ren}}^{k})$ at the pixel level using a combination of $L_{1}$ and D-SSIM terms, following the photometric error formulation in 3DGS [kerbl20233d] (Eq. 1) with $\lambda=0.2$ following 3DGS [kerbl20233d]. We normalize $C_{\text{pixel}}^{k}$ to $[0,1]$.
@@
Feature-level change cues: To capture high-level semantic differences, we leverage the visual foundation model SAM2 [ravi2024sam] to extract dense feature maps $(f_{\text{inf}}^{k},f_{\text{ren}}^{k})$ for $(I_{\text{inf}}^{k},I_{\text{ren}}^{k})$. Each feature map is represented as $f\in\mathbb{R}^{\frac{h}{s}\times\frac{w}{s}\times d}$, where $h$ and $w$ denote the image height and width, $s$ is the patch size, and $d$ is the feature dimensionality. The feature-level change cues $C_{\text{feature}}^{k}$ are computed as the absolute difference between the two feature maps:
@@
This formulation leverages the complementary strengths of pixel- and feature-level cues while avoiding the loss of change information. Pixel-level cues effectively capture fine-grained appearance differences, such as color variations between semantically similar objects, but tend to be more sensitive to distractor changes caused by shadows, reflections, or illumination shifts. Feature-level cues are more robust to these distractors yet may struggle to detect subtle differences within semantically similar regions. MV3DCD [galappaththige2025multi] relies on hard thresholding, which can disregard subtle but relevant changes that fall below predefined thresholds. Moreover, since MV3DCD fuses its structure and feature-aware masks through intersection, it may further lose valid change information not simultaneously captured by both masks. When combined with our novel self-supervised fusion loss (Sec. 3.4), the proposed formulation jointly integrates information across all observed viewpoints, effectively suppressing inconsistent distractors while maintaining sensitivity to meaningful changes.
@@ 3.4 Inferring Change Masks
MV3DCD [galappaththige2025multi] first enforced multi-view consistency for SCD using 3DGS [kerbl20233d]. We depart from its hard-thresholded heuristic fusion and introduce a novel self-supervised loss that jointly infers multi-view consistent change masks from all observed cues at test time.
After the reference scene representation $\mathcal{R}_{\text{ref}}$ is constructed (Sec. 3.1), we initialize the change representation $\mathcal{R}_{\text{change}}$ from $\mathcal{R}_{\text{ref}}$ by discarding all color parameters and introducing a learnable change parameter $c$ for each primitive.
@@ 3.5 Scene Representation Update
We then discard $c$ from each primitive in $\mathcal{R}_{\text{change}}$ and introduce view-dependent appearance modeled via spherical harmonics [kerbl20233d]. To only reconstruct changed regions, we mask the inference images using the refined change masks as $\hat{I}^{k}_{\text{inf}}=I^{k}_{\text{inf}}\odot M^{k}_{\text{refined}}$. $\hat{I}^{k}_{\text{inf}}$ guides the reconstruction of changed regions $\mathcal{R}_{\text{change}}^{*}$ following the standard 3DGS [kerbl20233d] optimization pipeline. This disentangled reconstruction is highly efficient and requires a fraction of the primitives compared to modeling the entire scene, thereby accelerating rendering and avoiding redundant computations in unchanged regions. Notably, this selective reconstruction achieves rendering speeds exceeding 400 FPS, substantially reducing overall optimization time.
Next, we fuse $\mathcal{R}^{*}_{\text{ref}},\mathcal{R}_{\text{change}}^{*}]$ to form the inference scene $\mathcal{R}_{\text{inf}}$, where $\mathcal{R}^{*}_{\text{ref}}$ denotes $\mathcal{R}_{\text{ref}}$ excluding primitives that contribute to changed pixels. A fast global optimization is then performed for 3k iterations, guided by $\mathcal{I}_{\text{inf}}$. We restrict the adaptive density control [kerbl20233d] only to the primitives contributing to changed pixels in at least one view to avoid unnecessary densification in unchanged regions.
@@ 4 Experiments
Datasets: We evaluate our method on PASLCD [galappaththige2025multi] for SCD. PASLCD comprises 10 room-scale (i.e., a cantina) indoor and outdoor scenes captured under similar and varying lighting conditions (20 instances in total). It features both surface-level appearance and object-level geometric changes, along with numerous distractors such as shadows, reflections, and illumination shifts, making it highly challenging multi-view dataset for SCD. Importantly, PASLCD captures scenes from unconstrained, independently traversed camera trajectories, closely reflecting real-world autonomous operation. For the scene representation update, we evaluate on PASLCD [galappaththige2025multi] and CL-Splats [ackermann2025clsplats]. CL-Splats consists of five small-scale (i.e., tabletop) scenes, each featuring a single object-level change.
Table 1: Quantitative results for SCD on PASLCD [galappaththige2025multi] averaged over all 20 instances. LF: Label-Free, PA: Pose-Agnostic, MV: Multi-View consistency for change detection, ON: Online. We additionally report the total runtime, including pose estimation and reference reconstruction, for offline methods, and the operating frame rate (FPS) for online methods. Our method achieves the best performance in both settings, even outperforming all existing offline methods while operating online.
@@
Baselines and Metrics: We conduct a comprehensive evaluation using the best-performing baselines [lin2025robust, cyws2d, kim2025towards, cho2025zero, lu20253dgs, galappaththige2025multi, park2021changesim, zhou2023pad, kruse2024splatpose, liu2024splatpose+]. For pairwise methods [lin2025robust, cyws2d, kim2025towards, cho2025zero, lu20253dgs] evaluated in the offline setting, we render identical viewpoints using vanilla 3DGS [kerbl20233d] for a fair comparison, although this substantially simplifies the task by removing viewpoint inconsistencies. For the online setting, we construct two additional baselines by integrating ChangeSim’s [park2021changesim] frame matching with the best-performing pairwise methods [cyws2d, kim2025towards]. We provide ground-truth poses for ChangeSim’s frame retrieval module to ensure a fair evaluation, as PASLCD lacks depths for ChangeSim’s off-the-shelf RGB-D SLAM system. We use model checkpoints provided by the authors for the supervised methods [lin2025robust, cyws2d]. Following standard practice in SCD [galappaththige2025multi, lin2025robust, alcantarilla2018street, sakurada2020weakly], we report mean intersection over union (mIoU) and F1 score.
For efficient scene representation update, we adopt 3DGS [kerbl20233d] as our primary baseline and further include two fast 3DGS variants [HansonSpeedy, hoellein_2025_3dgslm] to demonstrate the efficiency of our approach. We also evaluate CLNeRF [cai2023clnerf] (InstantNGP [muller2022instant] based) among publicly available continual learning methods. Following standard evaluation protocols [cai2023clnerf, ackermann2025clsplats, zeng2025gaussianupdate, wu2023cl, kerbl20233d], we report PSNR, SSIM, and LPIPS for novel views after scene update, along with runtimes.
@@ 4.1 Experiments on Scene Change Detection
Table 2: Runtime analysis of each module in our online SCD pipeline, measured in milliseconds per frame on PASLCD [galappaththige2025multi]. Most of the computation time is spent on multi-view change information fusion, while other modules are lightweight.
@@
Figure 5: Additional qualitative comparison with MV3DCD [galappaththige2025multi]. Our rendered change masks align more closely with the ground truth, capturing subtle structural and appearance changes that MV3DCD often misses. In contrast to MV3DCD, our method produces fewer spurious detections and demonstrates strong robustness to distractor variations across both indoor and outdoor environments.
@@
Runtime Analysis: Table 2 summarizes the runtime breakdown of our online SCD pipeline for a single inference frame $I^{k}_{\text{inf}}$. Approximately 33% of the total time is spent on pose estimation, aligning $I^{k}_{\text{inf}}$ to the coordinate system of $\mathcal{R}_{\text{ref}}$. Change detection accounts for 67% of the overall runtime, of which the majority (94%) is attributed to multi-view change cue fusion spent on optimizing $\mathcal{R}_{\text{change}}$.
@@
Qualitative Results: We present qualitative comparisons against our closest competitor, MV3DCD [galappaththige2025multi], in Figs. 2 and 5. Our method demonstrates superior change localization and robustness across diverse real-world scenes. Unlike MV3DCD, our approach produces cleaner and more spatially coherent change masks with significantly fewer false negatives. It effectively captures subtle appearance and geometric changes that MV3DCD often overlooks, while suppressing false positives caused by distractors. These results highlight that our self-supervised fusion of pixel- and feature-level cues combined with multi-view consistency enables accurate, fine-grained change detection.
Table 3: Ablation study of our SCD approach on PASLCD [galappaththige2025multi]. Performance benefits from every component.
@@
Ours (With [galappaththige2025multi]’s Thresholding & Heuristic Fusion) 0.350 0.495
@@
3DGS [kerbl20233d] 22.21 0.7558 0.2426 550 30.31 0.9319 0.1178 364
3DGS-LM [hoellein_2025_3dgslm] 22.26 0.7562 0.2422 340 29.95 0.9322 0.1177 275
SpeedySplats [HansonSpeedy] 22.25 0.7603 0.2618 399 29.89 0.9349 0.1290 312
CLNeRF [cai2023clnerf] 22.27 0.6239 0.3907 451 26.29 0.7867 0.2235 301
@@
Figure 6: Qualitative comparison of rendered views from the updated representation with CLNeRF [cai2023clnerf] and 3DGS [kerbl20233d] (from scratch). Our method more accurately reconstructs changed regions (red boxes) while reusing primitives from $\mathcal{R}_{\text{ref}}$ to preserve high fidelity in unchanged areas (yellow boxes), compared to naïve reconstruction at each time.
Ablation Analysis: We conduct an ablation study on PASLCD [galappaththige2025multi] (Table 3). Removing either the $L_{1}$ or $L_{\text{D-SSIM}}$ term in $C_{\text{pixel}}$ noticeably degrades performance, with the former providing a relatively stronger supervisory signal for $L_{\text{SSF}}$. When using either $C_{\text{pixel}}$ or $C_{\text{feature}}$, the model fails to converge, indicating that neither modality alone sufficiently guides $L_{\text{SSF}}$. Removing the regularization term collapses training into the trivial solution $\tilde{M}^{k}=1$ (disscussed in sec. 3.4). Using MV3DCD’s [galappaththige2025multi] hard thresholding and intersection heuristic instead of $L_{\text{SSF}}$ degrades performance.
Table 5: Analysis on scene update component on PASLCD [galappaththige2025multi]. Runtime for change detection and refinement is excluded. GO: Global Optimization, SR: Selective Reconstruction.
@@ 4.2 Experiments on 3DGS Representation Update
Qualitative Results. We present qualitative comparisons with 3DGS [kerbl20233d] and CLNeRF [cai2023clnerf] in Fig. 6. Our method more accurately reconstructs the changed regions compared to CLNeRF, while also achieving higher visual fidelity in unchanged areas than 3DGS (built from scratch), owing to the effective reuse of primitives from the reference scene.
@@ 5 Conclusion
We proposed a novel approach to pose-agnostic SCD that detects change in an online manner with state-of-the-art performance, outperforming even the best offline methods. We introduce two key innovations to achieve this: an ultra-light PnP-based pose estimator and a self-supervised fusion loss for learning a multi-view consistent change representation. Additionally, we introduced a change-guided update strategy for 3DGS, reducing training overhead to seconds while retaining reconstruction fidelity. Future work may focus on developing richer complementary change cues, potentially improving both online and offline SCD performance.

### v3

@@ 3.3 Extracting Change Cues
Pixel-level change cues: We quantify the differences between $(I_{\text{inf}}^{k},I_{\text{ren}}^{k})$ at the pixel level using a combination of $L_{1}$ and D-SSIM terms, following the photometric error formulation in 3DGS [19] (Eq. 1) with $\lambda=0.2$ following 3DGS [19]. We normalize $C_{\text{pixel}}^{k}$ to $[0,1]$.
@@
Feature-level change cues: To capture high-level semantic differences, we leverage the visual foundation model SAM2-Tiny [37] to extract dense feature maps $(f_{\text{inf}}^{k},f_{\text{ren}}^{k})$ for $(I_{\text{inf}}^{k},I_{\text{ren}}^{k})$. Each feature map is represented as $f\in\mathbb{R}^{\frac{h}{s}\times\frac{w}{s}\times d}$, where $h$ and $w$ denote the image height and width, $s$ is the patch size, and $d$ is the feature dimensionality. The feature-level change cues $C_{\text{feature}}^{k}$ are computed as the absolute difference between the two feature maps:
@@
This formulation leverages the complementary strengths of pixel- and feature-level cues while avoiding the loss of change information. Pixel-level cues effectively capture fine-grained appearance differences, such as color variations between semantically similar objects, but tend to be more sensitive to distractor changes caused by shadows, reflections, or illumination shifts. Feature-level cues are more robust to these distractors yet may struggle to detect subtle differences within semantically similar regions. MV3DCD [13] relies on hard thresholding, which can disregard subtle but relevant changes that fall below predefined thresholds. Moreover, since MV3DCD fuses its structure and feature-aware masks through intersection, it may further lose valid change information not simultaneously captured by both masks. When combined with our novel self-supervised fusion loss (Sec. 3.4), the proposed formulation jointly integrates information across all observed viewpoints, effectively suppressing inconsistent distractors while maintaining sensitivity to meaningful changes.
@@ 3.4 Inferring Change Masks
MV3DCD [13] first enforced multi-view consistency for SCD using 3DGS [19]. We depart from its hard-thresholded heuristic fusion and introduce a novel self-supervised loss that jointly infers multi-view consistent change masks from all observed cues at test time.
After the reference scene $\mathcal{R}_{\text{ref}}$ is constructed (Sec. 3.1), we initialize the change representation $\mathcal{R}_{\text{change}}$ from $\mathcal{R}_{\text{ref}}$ by discarding all color parameters and introducing a learnable change parameter $c$ [13] for each primitive.
@@ 3.5 Scene Representation Update
We then discard $c$ from each primitive in $\mathcal{R}_{\text{change}}$ and introduce view-dependent appearance modeled via spherical harmonics [19]. To only reconstruct changed regions, we mask the inference images using the refined change masks as $\hat{I}^{k}_{\text{inf}}=I^{k}_{\text{inf}}\odot M^{k}_{\text{refined}}$. $\hat{I}^{k}_{\text{inf}}$ guides the reconstruction of changed regions $\mathcal{R}_{\text{change}}^{*}$ following the standard 3DGS [19] optimization pipeline. This disentangled reconstruction is highly efficient and requires a fraction of the primitives compared to modeling the entire scene, thereby accelerating rendering and avoiding redundant computations in unchanged regions. Notably, this selective reconstruction achieves rendering speeds exceeding 400 FPS, substantially reducing overall optimization time.
Next, we fuse $\mathcal{R}^{*}_{\text{ref}},\mathcal{R}_{\text{change}}^{*}]$ to form the inference scene $\mathcal{R}_{\text{inf}}$, where $\mathcal{R}^{*}_{\text{ref}}$ denotes $\mathcal{R}_{\text{ref}}$ excluding primitives that contribute to changed pixels. A fast global optimization is then performed, guided by $\mathcal{I}_{\text{inf}}$. We restrict the adaptive density control [19] only to the primitives contributing to changed pixels in at least one view to avoid unnecessary densification in unchanged regions.
@@ 4 Experiments
Datasets: We evaluate our method on PASLCD [13] for SCD. PASLCD comprises 10 room-scale (i.e., a cantina) indoor and outdoor scenes captured under similar and varying lighting conditions. It features both surface-level appearance and object-level geometric changes, along with numerous distractors such as shadows, reflections, and illumination shifts, making it a highly challenging multi-view dataset for SCD. Importantly, PASLCD captures scenes from unconstrained, independently traversed camera trajectories, closely reflecting real-world autonomous operation. For the scene representation update, we evaluate on PASLCD [13] and CL-Splats [1]. CL-Splats consists of five small-scale (i.e., tabletop) scenes, each featuring a single object-level change.
Table 1: Quantitative results for SCD on PASLCD [13] averaged over all 20 instances. LF: Label-Free, PA: Pose-Agnostic, MV: Multi-View consistency for change detection, ON: Online. We additionally report the total runtime, including pose estimation and reference reconstruction, for offline methods, and the operating frame rate (FPS) for online methods. Our method achieves the best performance in both settings, even outperforming all existing offline methods while operating online.
@@
Baselines and Metrics: We conduct a comprehensive evaluation using the best-performing baselines [28, 39, 20, 8, 30, 13, 34, 52, 22, 29]. For pairwise methods [28, 39, 20, 8, 30] evaluated in the offline setting, we render identical viewpoints using vanilla 3DGS [19] for a fair comparison, although this substantially simplifies the task by removing viewpoint inconsistencies. For the online setting, we construct two additional baselines by integrating ChangeSim’s [34] frame matching with the best-performing pairwise methods [39, 20]. We provide ground-truth poses for ChangeSim’s frame retrieval module to ensure a fair evaluation, as PASLCD lacks depths for ChangeSim’s off-the-shelf RGB-D SLAM system. We use model checkpoints provided by the authors for the supervised methods [28, 39]. Following standard practice in SCD [13, 28, 2, 41], we report mean intersection over union (mIoU) and F1 score computed for change pixels in the ground-truth mask.
For efficient scene representation update, we adopt 3DGS and fast variants [15, 16, 19] as our baselines. We also evaluate CLNeRF [5] among publicly available continual learning methods. Following standard evaluation protocols [5, 1, 51, 49, 19], we report PSNR, SSIM, and LPIPS for novel views after scene update, along with runtimes.
@@ 4.1 Experiments on Scene Change Detection
Offline SCD Results: Table 1 presents an extensive comparison against SOTA methods across all SCD settings on PASLCD [13]. In the offline setting, we follow the protocol of MV3DCD [13] by optimizing $\mathcal{R}_{\text{change}}$ using our $L_{\text{SSF}}$ with access to all inference views jointly for 3k iterations.
@@
Table 2: Runtime analysis of each module in our online SCD pipeline, measured in milliseconds per frame on PASLCD [13]. Most of the computation time is spent on multi-view change information fusion, while other modules are lightweight.
@@
Figure 5: Additional qualitative comparison with MV3DCD [13]. Our rendered change masks align more closely with the ground truth, capturing subtle structural and appearance changes that MV3DCD often misses. In contrast to MV3DCD, our method produces fewer spurious detections and demonstrates strong robustness to distractor variations across both indoor and outdoor environments.
@@
Runtime Analysis: Table 2 summarizes the runtime breakdown (asynchronous overhead) of our online SCD pipeline for a single inference frame $I^{k}_{\text{inf}}$. Approximately 33% of the total time is spent on pose estimation, aligning $I^{k}_{\text{inf}}$ to the coordinate system of $\mathcal{R}_{\text{ref}}$. SCD accounts for 67% of the overall runtime, of which the majority (94%) is attributed to multi-view change cue fusion spent on optimizing $\mathcal{R}_{\text{change}}$.
@@
Qualitative Results: We present qualitative comparisons against our closest competitor, MV3DCD [13], in Figs. 2 and 5. Our method demonstrates superior change localization and robustness across diverse real-world scenes. Unlike MV3DCD, our approach produces cleaner and more spatially coherent change masks with significantly fewer false negatives. It effectively captures subtle appearance and geometric changes that MV3DCD often overlooks, while suppressing false positives caused by distractors. These results highlight that our self-supervised fusion of pixel- and feature-level cues combined with multi-view consistency enables accurate, fine-grained change detection.
Table 3: Ablation study of our SCD approach on PASLCD [13]. Performance benefits from every component.
@@
Ours (With [13]’s Thresholding & Heuristic Fusion) 0.350 0.495
@@
3DGS [19] 22.21 0.7558 0.2426 550 30.31 0.9319 0.1178 364
3DGS-LM [16] 22.26 0.7562 0.2422 340 29.95 0.9322 0.1177 275
SpeedySplats [15] 22.25 0.7603 0.2618 399 29.89 0.9349 0.1290 312
CLNeRF [5] 22.27 0.6239 0.3907 451 26.29 0.7867 0.2235 301
@@
Figure 6: Qualitative comparison of rendered views from the updated representation with CLNeRF [5] and 3DGS [19] (from scratch). Our method more accurately reconstructs changed regions (red boxes) while reusing primitives from $\mathcal{R}_{\text{ref}}$ to preserve high fidelity in unchanged areas (yellow boxes), compared to naïve reconstruction at each time.
Ablation Analysis: We conduct an ablation study on PASLCD [13] (Table 3). Removing either the $L_{1}$ or $L_{\text{D-SSIM}}$ term in $C_{\text{pixel}}$ noticeably degrades performance, with the former providing a relatively stronger supervisory signal for $L_{\text{SSF}}$. When using either $C_{\text{pixel}}$ or $C_{\text{feature}}$, the model fails to converge, indicating that neither modality alone sufficiently guides $L_{\text{SSF}}$. Removing the regularization term collapses training into the trivial solution $\tilde{M}^{k}=1$ (disscussed in sec. 3.4). Using MV3DCD’s [13] hard thresholding and intersection heuristic instead of $L_{\text{SSF}}$ degrades performance.
Table 5: Analysis on scene update component on PASLCD [13]. Runtime for change detection and refinement is excluded. GO: Global Optimization, SR: Selective Reconstruction.
@@ 4.2 Experiments on 3DGS Representation Update
Qualitative Results: We present qualitative comparisons with 3DGS [19] and CLNeRF [5] in Fig. 6. Our method more accurately reconstructs the changed regions compared to CLNeRF, while also achieving higher visual fidelity in unchanged areas than 3DGS (built from scratch), owing to the effective reuse of primitives from the reference scene.
@@ 5 Conclusion
We proposed a novel approach to pose-agnostic SCD that detects change in an online manner with SOTA performance, outperforming even the best offline methods. We introduce two key algorithmic and system-level innovations to achieve this: an ultra-light PnP-based pose estimator and a self-supervised fusion loss for learning a multi-view consistent change representation. Additionally, we introduced a change-guided update strategy for 3DGS, reducing training overhead to seconds while retaining reconstruction fidelity. Future work may focus on developing richer complementary change cues, potentially improving both online and offline SCD performance.